Circular No. 5843

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 PB J. V. Scotti, Lunar and Planetary Laboratory, reports his discovery of

a fast-moving asteroidal object with the Spacewatch telescope, Kitt Peak:

1993 UT  α_{2000}  δ_{2000}  V
Aug. 13.19268  21^{h}13^{m}24.^{s}40  −10^{o}09′10′′.9  18.0
13.22507  21 13 18.68  −10 08 58.4  17.9
13.25864  21 13 12.78  −10 08 44.7  18.0
14.24796  21 10 22.56  −10 02 12.0  17.8
14.25743  21 10 21.00  −10 02 07.8
14.26667  21 10 19.32  −10 02 04.2  18.0
15.33552  21 07 18.46  − 9 55 04.5  17.9
15.34303  21 07 17.18  − 9 55 01.5  17.9
15.35614  21 07 14.92  − 9 54 56.5  17.9

Very preliminary orbital elements from the above observations:

T = 1993 Feb. 22.394 TT  ω = 211.^{o}738
e = 0.61189  Ω = 316.178 }2000.0
q = 0.55689 AU  i = 42.664

a = 1.43486 AU n^{o} = 0.573442 P = 1.72 years

1993 TT α_{2000} δ_{2000} Δ r ϵ β V
Aug. 11 21^{h}19.^{m}85 −10^{o}23.^{′}8 0.936 1.947 175.^{o}0 2.^{o}6 17.6
16 21 05.47 − 9 50.7 0.968 1.973 170.2  5.0 17.8
21 20 52.39 − 9 19.4 1.009 1.998 162.9  8.5 18.1
26 20 40.87 − 8 50.4 1.057 2.022 155.7 11.9 18.4
31 20 30.99 − 8 23.8 1.113 2.044 148.7 14.9 18.6
Sept. 5 20 22.78 − 7 59.7 1.175 2.066 142.1 17.4 18.8
10 20 16.17 − 7 37.8 1.243 2.087 135.8 19.6 19.1

PERSEID METEORS 1993 Further to IAUC 5841, M. Gyssens, International Meteor Organization,

informs us that the increased Perseid activity clearly extended beyond Aug. 12.5 UT and that after that time there could be confusion with the expected peak of the ‘normal’ Perseid shower. A report relayed from E. Mashal, Tel-Aviv University, indicates that at the Wise Observatory Perseids were seen under perfect conditions at an observed hourly rate of 720 for ∼ 30 min around Aug. 12.04 UT.

1993 August 16 Brian G. Marsden